### v3

# NBBOX: Noisy Bounding Box Improves Remote Sensing Object Detection

Abstract
Data augmentation has shown significant advancements in computer vision to improve model performance over the years, particularly in scenarios with limited and insufficient data. Currently, most studies focus on adjusting the image or its features to expand the size, quality, and variety of samples during training in various tasks including object detection. However, we argue that it is necessary to investigate bounding box transformations as a data augmentation technique rather than image-level transformations, especially in aerial imagery due to potentially inconsistent bounding box annotations. Hence, this letter presents a thorough investigation of bounding box transformation in terms of scaling, rotation, and translation for remote sensing object detection. We call this augmentation strategy NBBOX (Noise Injection into Bounding Box). We conduct extensive experiments on DOTA and DIOR-R, both well-known datasets that include a variety of rotated generic objects in aerial images. Experimental results show that our approach significantly improves remote sensing object detection without whistles and bells and it is more time-efficient than other state-of-the-art augmentation strategies.
Index Terms:
Noisy bounding box, oriented bounding box, remote sensing object detection, data augmentation.
I Introduction
In recent years, remote sensing object detection has achieved remarkable success primarily due to the advancements in deep learning. Particularly, modern deep learning architectures such as convolutional neural networks and transformers have significantly enhanced the capabilities of object detection in Earth vision (e.g. [1, 2]). These models have millions or billions of parameters to learn, requiring tremendous training data to avoid the over-fitting problem. Compared to natural scenes (e.g. ImageNet [3] or MS-COCO [4]), datasets for overhead imagery usually lack diversity and quantity due to the expense of data collection. To overcome this limitation, data augmentation plays a significant role as an implicit regularization strategy for model learning.
Existing visual data augmentation methods can be mainly divided into three categories: (a) image manipulation (e.g. [5, 6, 7, 8]), (b) feature transformation (e.g. [9, 10]), and (c) generative image synthesis (e.g. [11, 12, 13]). For (a), geometric transformations (such as flipping and shearing), sharpness transformations, noise distribution, etc are considered the simplest augmentations. More advanced strategies in image manipulation contain image erasing and image mix. The image erasing like [5, 6] randomly selects a sub-region in an image and deletes its contents, while the image mix like [7, 8] combines two or more images (or patches) as a single image during training. Meanwhile, (b) aims to directly manipulate features in the latent space rather than that raw input image. For example, [9] applies augmentations to the extracted features and encourages the model to be consistent for both the original and augmented features. [10] leverages the synergy between feature normalization and data augmentation to improve the robustness and accuracy of models. Unfortunately, both are hard to directly apply to rotated object detection. Finally, (c) produces samples with generative models based on GAN [11] or Diffusion [13]. Such methods can generate high-quality (for GAN) or high-fidelity (for Diffusion) samples, but these require significant computational resources.
While most prior work for data augmentation concentrates on adjusting the image or its features, only few have focused on the deformation of bounding boxes for object detection. However, [14] discovers the presence of low-confidence annotations in current overhead object detection datasets: there is a mismatch between the minimum enclosing box (i.e. optimal) and the actual annotation. Thus, we argue that investigating bounding box transformation is necessary to boost the model performance for remote sensing object detection, by enabling robust training in potentially inaccurate bounding boxes.
In this letter, we propose a simple but efficient data augmentation method named NBBOX (Noise Injection into Bounding Box) for remote sensing object detection. Intuitively, our method adds noise to bounding boxes with geometric transformations during training (not for test/deployment). Particularly, this work thoroughly investigates bounding box transformation in terms of scaling, rotation, and translation. To demonstrate the effectiveness of our approach, extensive experiments are conducted on DOTA [15] and DIOR-R [16], both datasets that contain rotated and densely-placed objects.
Figure 1: Examples of the proposed data augmentation method named NBBOX for remote sensing object detection.
Figure 2: Comparison between provided bounding box labels and minimum enclosing rectangles for objects on DOTA and DIOR-R.
II Methodology
In this section, we present a simple, but efficient data augmentation strategy namely NBBOX (Noise Injection into Bounding Box) for remote sensing object detection. As seen in Fig. 2, there is often a discrepancy between the actual annotation and the minimum enclosing box (i.e. optimal localization label) on the existing remote sensing datasets, which might cause difficulties in model training due to inaccurate and low-consistent bounding boxes. To enhance the model’s robustness against such implicit box-level noise, our augmentation scheme randomly adds noise to oriented bounding boxes with the three simplest geometric transformations: scaling, rotation, and translation during training, as depicted in Fig. 1.
II-A Background
Training models under label noise is challenging and requires careful consideration to learn resilient and generalized models. Most research attention has been given to the classification task, but recently a few have dedicated to object detection (e.g. [17, 18, 19, 20] for natural scenes and [21, 22, 23] for remote sensing imagery).
For natural images, [17] first analyzes the effects of various label noise types on object detection and introduces a per-object co-teaching framework to address the noise label issue. [18] suggests a learning strategy of alternating between rectifying noise and training the model to handle label noise concerning category and bounding box. [19] proposes to address label noise by adopting meta-learning via exploiting a few clean labels. Previous studies consider classification and localization simultaneously, whereas [20] employs multiple instance learning to deal solely with inaccurate bounding boxes. We believe that it is reasonable and practical to pursue this research direction as imprecise bounding box labels are more prevalent than noisy category labels in practice.
For remote sensing images, a few work has been proposed to address noisy labels in object detection, inspired by research for natural scenes. [21] designs two kinds of loss functions to mitigate mislabels for both classification and localization. On the other hand, [22], [23] focus on correcting incomplete annotations such as missing bounding boxes due to misalignment of GPS sensors. Unlike prior work, the main goal of our work is to resolve localization errors inherent in the existing bounding box labels. In other words, we consider a generic object detection scenario, where missing bounding boxes or incorrect class labels of the ground truth are uncommon. Instead, we assume that the bounding box labels may not be optimal due to the annotators’ mistakes or dilemma situations such as how to rotate a bounding box for a circular object on overhead imagery during the annotation process. In particular, we study a data augmentation strategy based on bounding box transformation for more robust localization.
II-B Overview of the proposed approach
The detail of NBBOX is presented in Algorithm 1. In our algorithm, there exist two types of input: $\mathbf{I}$ is the input image and $\mathbf{L}=\{(\mathbf{B}_{i},\mathbf{c}_{i})\}^{N}_{i=1}$ is the corresponding labels of $N$ objects for object detection, where $\mathbf{B}_{i}=({x}_{\text{c}},{y}_{\text{c}},w,h,\theta)$ and $\mathbf{c}_{i}$ correspond to the bounding box label and the one-hot encoded category label for $i$-th object, respectively. $({x}_{\text{c}},{y}_{\text{c}})$ is the center point, while $w$ and $h$ represent the width and height of the bounding box. $\theta$ denotes the rotated angle of the box.
Moreover, there are three kinds of hyper-parameters to tune. $\mathbf{s}=({s}_{\text{min}},{s}_{\text{max}})$, $\mathbf{r}=({r}_{\text{min}},{r}_{\text{max}})$, and $\mathbf{t}=({t}_{\text{min}},{t}_{\text{max}})$ determine the ranges of translation, scaling, and rotation, respectively, where $\forall{k},{s}_{k}\in\mathbb{R}\wedge{r}_{k}\in\mathbb{R}\wedge{t}_{k}\in\mathbb{Z}$. Here, $\mathbb{Z}$ is the set of all integers whereas $\mathbb{R}$ is for all real numbers.
${\text{rand}}_{\text{int}}(a,c)\sim{P}_{\text{int}}$ selects and returns any specific integer $c$ such that $a\leq b\leq c$. ${\text{rand}}_{\text{float}}(a,c)\sim{P}_{\text{float}}$ generates any real number value within the closed interval $[a,c)$. We simply choose the uniform distributions for all random functions used in our algorithm. In other words, ${P}(b;a,c)=\frac{1}{c-a+1}$ and ${P}(b;a,c)=\frac{1}{c-a}$ are used for the probability mass (discrete) function ${P}_{\text{int}}$ and the probability density (continuous) function ${P}_{\text{float}}$, respectively. One might question why $\mathbf{t}$ has a domain of integer values while $\mathbf{s}$ and $\mathbf{r}$ have domains of real numbers. This is because translation operations are conducted on a pixel basis, while scaling and rotation factors can be the real number.
Assume that a single image $\mathbf{I}$ is given with labels $\mathbf{L}$ to our algorithm. The ultimate goal is to perform bounding box transformations during training as a data augmentation strategy. To clarify, we randomly add noise to each bounding box $\mathbf{B}_{i}$ in $\mathbf{L}$ for each training epoch. In Algorithm 1, for each box $\mathbf{B}_{i}$, three kinds of transformations are applied in sequence (scaling: line 4-5, rotation: line 7-8, translation: line 10-11). Note that it is possible to activate only a subset of these sub-operations during the execution of NBBOX. For instance, you can simply choose not to use ‘rotation’ for NBBOX if needed.
1
Input : Input image $\mathbf{I}$, Labels $\mathbf{L}=\{(\mathbf{B}_{i},\mathbf{c}_{i})\}^{N}_{i=1}$
Parameters : Scaling range $\mathbf{s}=({s}_{\text{min}},{s}_{\text{max}})$,
Rotation range $\mathbf{r}=({r}_{\text{min}},{r}_{\text{max}})$,
Translation range $\mathbf{t}=({t}_{\text{min}},{t}_{\text{max}})$
Output : Input image $\mathbf{I}$, Updated labels $\mathbf{L}^{*}$
2 for $i\leftarrow 1$ to $N$ do
3       ${x}_{\text{c}},{y}_{\text{c}},w,h,\theta$leftarrow $\mathbf{B}_{i}$;
4       // (1) BBOX Scaling ;
5       $s_{w},s_{h}\leftarrow\text{rand}_{\text{float}}({s}_{\text{min}},{s}_{\text{max}}),\text{rand}_{\text{float}}({s}_{\text{min}},{s}_{\text{max}})$;
6       $w^{*},h^{*}\leftarrow w\times s_{w},h\times s_{h}$;
7       // (2) BBOX Rotation ;
8       $r_{\theta}\leftarrow\text{rand}_{\text{float}}({r}_{\text{min}},{r}_{\text{max}})$;
9       $\theta^{*}\leftarrow\theta+r_{\theta}$;
10       // (3) BBOX Translation ;
11       $t_{x},t_{y}\leftarrow\text{rand}_{\text{int}}({t}_{\text{min}},{t}_{\text{max}}),\text{rand}_{\text{int}}({t}_{\text{min}},{t}_{\text{max}})$;
12       ${x}^{*}_{\text{c}},{y}^{*}_{\text{c}}\leftarrow{x}_{\text{c}}+t_{x},{y}_{\text{c}}+t_{y}$;
13       // Update ;
14       ${\mathbf{B}}^{*}_{i}\leftarrow{x}^{*}_{\text{c}},{y}^{*}_{\text{c}},w^{*},h^{*},\theta^{*}$;
15
[…]
Algorithm 1 Procedure of NBBOX.
II-C Advanced options for translation and scaling
As demonstrated in Algorithm 1, the bounding box may be translated by different values along both the x-axis and y-axis. Besides, the aspect ratio may not be maintained for scaling operations. This phenomenon would not be recommended for training on certain data scenarios. For instance, it would not be desirable to modify the aspect ratio of bounding boxes in an extreme situation where most objects are expected to have a square size. Hence, we alleviate this by providing additional parameters ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ in our algorithm. Here, ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ are boolean values to be set by users. One can effortlessly modify NBBOX with ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ as follows.
For scaling, line 4 should be replaced with:
$\begin{split}&\alpha,\beta\leftarrow{\text{rand}}_{\text{float}}({s}_{\text{min}},{s}_{\text{max}}),{\text{rand}}_{\text{float}}({s}_{\text{min}},{s}_{\text{max}});\\ &s_{w}\leftarrow\alpha\text{ if }{\text{bool}}_{s}\text{ is True, otherwise }\alpha;\\ &s_{h}\leftarrow\alpha\text{ if }{\text{bool}}_{s}\text{ is True, otherwise }\beta,\\ \end{split}$ (1)
where ${\text{bool}}_{s}$ represents whether the bounding box is proportionally resized to keep the width and height ratio unchanged.
Likewise, for translation, line 10 should be replaced with:
$\begin{split}&\alpha,\beta\leftarrow{\text{rand}}_{\text{int}}({t}_{\text{min}},{t}_{\text{max}}),{\text{rand}}_{\text{int}}({t}_{\text{min}},{t}_{\text{max}});\\ &t_{x}\leftarrow\alpha\text{ if }{\text{bool}}_{t}\text{ is True, otherwise }\alpha;\\ &t_{y}\leftarrow\alpha\text{ if }{\text{bool}}_{t}\text{ is True, otherwise }\beta,\\ \end{split}$ (2)
where ${\text{bool}}_{t}$ denotes whether the center point is shifted equally along both the $x$ and $y$ directions.
II-D Scale-aware noise injection into bounding box
Tiny objects often occupy only a few pixels in an image, making their bounding boxes highly sensitive to even minor perturbations. For tiny objects, such transformations can lead to significant discrepancies between the transformed bounding box and the actual object, resulting in huge misalignment [24]. To fill this gap, we propose enhancing our algorithm by replacing line 2 in Algorithm 1 with:
$\begin{split}&{x}_{\text{c}},{y}_{\text{c}},w,h,\theta\leftarrow\mathbf{B}_{i};\\ &\text{if }w\leq\gamma\text{ or }h\leq\gamma:\text{continue},\\ \end{split}$ (3)
where $\gamma\in\mathbb{Z}$ is a hyper-parameter to tune. For example, $\gamma=0$ means applying random transformations to all bounding boxes, as in our original algorithm. In contrast, when $\gamma$ is a positive integer, it indicates that bounding boxes with a width or height smaller than $\gamma$ will remain unaltered. Hence, the proper value of $\gamma$ helps preserve the integrity of tiny objects, ensuring that their spatial properties are not compromised by transformations that may lead to misalignment or loss of information in remote sensing imagery. Note that our Python implementation with Numpy slightly deviates from our pseudo-code for efficient computation (See supp. material).
III Experiments
In this section, we show our data augmentation strategy NBBOX improves the performance of remote sensing object detection. We first briefly introduce the experimental settings and implementation details. We then provide test results and analysis of experiments with the proposed method.
III-A Experimental setup
Detection models. Traditionally, deep learning-based object detection models are typically separated into one-stage and two-stage methods. Nowadays, the distinction between anchor-based and anchor-free methodologies tends to be more emphasized when classifying detector models. This work adopts rotated Faster R-CNN [25] and FCOS [26], representative models in anchor-based two-stage and anchor-free one-stage detection studies, respectively. For all models, either ImageNet-pretrained ResNet-50 [27], a representative CNN model, or Swin-T [28], a representative Transformer model, is used for feature extraction. Additionally, FPN-1x [29] is incorporated as a neck to enhance multi-scale invariance.
Training details. We implement our hypothesis using PyTorch and MMRotate [30]. All the models are trained for 25 epochs. For a fair comparison with other approaches, all the experiments are resumed with weights pre-trained for five epochs without any augmentation per each {detector, data} combination. SGD is used for optimization with momentum of weight 0.9, weight decay 1e-4, learning rate 1e-3, and batch size 8. Learning rate warm-up is used over the first 500 iterations, where the learning rate linearly increases from zero to 33% of 5e-3. L2-norm gradient clipping is utilized with a maximum norm of 35. We apply the pre-processing operations ‘RResize’, ‘RRandomFlip’, ‘Normalize’, and ‘Pad’ for training, while ‘MultiScaleFlipAug’ is also used for testing.
Figure 3: Impact of scaling, rotation, and translation of bounding boxes on model performance: as we observe similar outcomes across different architectures and datasets, for brevity, we include only the result from Faster R-CNN with ResNet-50 trained on DIOR-R.
Evaluation metric. To objectively verify performance, all experiments are repeated five times, and the evaluation results are reported as ‘mean ± standard deviation’. For each experiment, we use mAP (mean Average Precision) at IoU($,\textbf{y}$ y)=0.5 (i.e. mAP@50) to measure the accuracy of the detected bounding boxes ($\hat{\textbf{y}}$) against the ground truth boxes (y) following [4]. This metric evaluates both the precision and recall of the model across different classes as:
$\text{mAP}=\frac{1}{C}\sum_{i=1}^{C}{\text{AP}}_{i},$ (4)
where $C$ is the number of categories and ${\text{AP}}_{i}$ is the average precision for category $i$. See [31] for detailed calculation.
III-B Datasets
This work conducts experiments on two widely recognized datasets for remote sensing object detection: DOTA (v1) [15] and DIOR-R [16]. Both have diverse real-world categories, high resolution, and detailed annotations. DOTA includes 15 object categories such as ‘car’, ‘ship’, and ‘harbor’. It contains 2,806 aerial images of varying resolutions with 188,282 oriented bounding box labels. Compared to DOTA, DIOR-R has 20 object categories including 192,512 instances over 23,463 images for rotated object detection. We conduct experiments and report results based on the standard convention like [32] in the remote sensing community for both datasets.
III-C Impact of scaling, rotation, and translation in NBBOX
This subsection studies the model performance when exposed to varying levels of scaling, rotation, and translation of bounding boxes. One fact is that too much deformation will adversely affect model training. In other words, to mitigate the bad effects of NBBOX, we assume that (i) ${s}_{\text{min}}$ and ${s}_{\text{max}}$ should be close to 1; (ii) ${r}_{\text{min}}$ and ${r}_{\text{max}}$ should be close to 0; (iii) ${t}_{\text{min}}$ should be close to -1 and ${t}_{\text{max}}$ should be close to 1. This reduces the difficulty of parameter tuning. To validate this assumption, we conduct experiments on DIOR-R with Faster R-CNN and FCOS, both employing ResNet-50 as the backbone (See Fig. 3 and supp. material for details). Here we consider combinations of the following parameters: ${s}_{\text{min}}\in$0.5, 0.8, 0.9, 0.95, 0.99}, ${s}_{\text{max}}\in$1, 1.01, 1.05, 1.1, 1.2, 1.5}, $\{{r}_{\text{min}},{r}_{\text{max}}\}\in$±0.01, ±0.1, ±0.5, ±1, ±2, ±3, ±4, ±5, ±10}, and $\{{t}_{\text{min}},{t}_{\text{max}}\}\in$±1, ±2, ±3, ±4, ±5, ±10, ±20, ±30}. As a result, we find parameters that support our hypothesis as:
  • ${s}_{\text{min}}\in$0.95, 1}, ${s}_{\text{max}}\in$1, 1.01};
  • $\{{r}_{\text{min}},{r}_{\text{max}}\}$ = {±0.01} & ${t}_{\text{min}}\in$-1, -2}, ${t}_{\text{max}}\in$1, 2}.
Unless otherwise specified, we use ${s}_{\text{min}}$ = 0.99, ${s}_{\text{max}}$ = 1.01, $\{{r}_{\text{min}},{r}_{\text{max}}\}$ = {±0.01}, and $\{{t}_{\text{min}},{t}_{\text{max}}\}$ = {±1} for remaining.
Isotropic scaling and translation. When scaling with preserving aspect ratio (i.e. ${\text{bool}}_{s}$=True), all models show a slight increase in mAP compared to when it is not (i.e. ${\text{bool}}_{s}$=False) as shown in Table VI. Besides, in the case of translation noise, the models perform slightly better when isotropic translation is applied (i.e. ${\text{bool}}_{t}$=True) compared to when it is not (i.e. ${\text{bool}}_{t}$=False) as indicated in Table VI. Thus, experimental results on DIOR-R in Table VI suggest that isotropic scaling and translation may be important. In the remaining experiments, we use ${\text{bool}}_{s}$=True and ${\text{bool}}_{t}$=True.
TABLE I: Average mAP under isotropic/non-isotropic scaling and
translation noises for each parameters
(${s}_{\text{min}}$, ${s}_{\text{max}}$, ${t}_{\text{min}}$, and ${t}_{\text{max}}$) on DIOR-R in NBBOX.
(With ResNet-50) Faster R-CNN FCOS
Scaling ${\text{bool}}_{s}$=True 60.12 (+0.05) 56.97 (+0.05)
${\text{bool}}_{s}$=False 60.07 (+0.00) 56.92 (+0.00)
Translation ${\text{bool}}_{t}$=True 60.37 (+1.08) 58.38 (+1.46)
${\text{bool}}_{t}$=False 59.29 (+0.00) 56.92 (+0.00)
TABLE II: Impact of combining scaling, rotation, and translation of bounding boxes on DIOR-R in NBBOX.
(Bold: best, underlined: second-best, wavy-underlined: third-best in this paper)
(With ResNet-50) Faster R-CNN FCOS
Scaling (Best) 62.79 ± 0.05 60.00 ± 0.99
Rotation (Best) 62.87 ± 0.29 59.68 ± 1.03
Translation (Best) 62.78 ± 0.26 59.93 ± 0.91
Scaling + Rotation 62.92 ± 0.24 60.13 ± 0.97
Scaling + Translation 62.89 ± 0.30 60.02 ± 1.07
Translation + Rotation 62.80 ± 0.22 59.89 ± 1.06
Scaling + Rotation + Translation 63.18 ± 0.14 60.25 ± 1.00
TABLE III: Impact of injecting scale-aware noise on DIOR-R in NBBOX.
(With ResNet-50) $\gamma$ = 0 $\gamma$ = 8 $\gamma$ = 16 $\gamma$ = 32 $\gamma$ = 64
Faster R-CNN 63.18 ± 0.14 63.20 ± 0.20 63.21 ± 0.18 63.17 ± 0.17 62.93 ± 0.27
FCOS 60.25 ± 1.00 60.24 ± 1.18 60.27 ± 1.06 59.99 ± 1.30 59.74 ± 1.24
Combining scaling, rotation, and translation. There is a reasonable expectation that integrating scaling, rotation, and translation could synergistically improve the model’s robustness and generalization. Hence, we further conduct experiments to validate the effectiveness of such combinations. Precisely, we study different combinations of each transformation (with parameters that show the best): pairwise combinations and the complete combination. Table VII indicates the effectiveness of the combined transformations in enhancing the model’s performance. Notably, the improvements are particularly evident in the scenario when we simultaneously use scaling, rotation, and translation during training.
III-D Impact of scale-aware noise injection in NBBOX
Following III-C, this subsection particularly focuses on a parameter study of $\gamma$, where $\gamma$ is varied over the set {0, 8, 16, 32, 64} under the combination of scaling, rotation, and translation. For this, we perform experiments on DIOR-R using Faster R-CNN and FCOS, both configured with ResNet-50 as the backbone as in III-C. As demonstrated in Table III, both detectors achieve their highest accuracies at $\gamma=16$, suggesting that bounding box perturbation for tiny objects (less than 16 $\times$ 16 pixels) might not be recommended. Besides, as $\gamma$ increases beyond 16, the performance degrades, indicating that moderate noise injection can enhance the model’s robustness. Hence, we use $\gamma$ = 16 for the remaining experiments.
III-E Comparison with state-of-the-art methods
In this subsection, we compare our method with easy-to-get data augmentation methods based on image manipulation (image/patch erasing and mix) as illustrated in Table IV. For a DOTA dataset, we use the same hyper-parameters for DIOR-R. Table IV shows that our NBBOX achieves a competitive performance among the state-of-the-art augmentation methods. Its performance can be further enhanced with other methods, such as RandRotate. Notably, obtaining satisfactory results on the DOTA dataset using the same parameters derived from DIOR-R demonstrates that the proposed method requires minimal parameter tuning. In addition, our method is faster than others: training Faster R-CNN with ResNet-50 on DIOR-R takes an average of 9.63 minutes per epoch for our method, compared to 21.24 minutes for MosaicMix. Due to the page limit, more results and analysis are in the supp. material.
TABLE IV: Comparison of state-of-the-art methods.
Backbone Data DIOR-R DOTA
Detectors Faster R-CNN FCOS Faster R-CNN FCOS
ResNet-50 Baseline 62.66 ± 0.17 59.75 ± 1.09 68.92 ± 0.44 67.00 ± 0.26
+ RandRotate [33] 63.10 ± 0.21 60.11 ± 1.12 70.02 ± 0.32 67.56 ± 0.49
+ RandShift [33] 62.87 ± 0.37 59.67 ± 1.05 69.01 ± 0.40 67.22 ± 0.59
+ CutOut [5] 62.91 ± 0.26 59.84 ± 0.99 69.13 ± 0.45 67.42 ± 0.37
+ AutoAug [34] 62.33 ± 0.31 59.24 ± 1.07 68.87 ± 0.32 67.45 ± 0.49
+ RandAug [35] 62.55 ± 0.12 59.94 ± 1.29 66.11 ± 0.24 65.08 ± 0.70
+ AugMix [36] 62.83 ± 0.11 59.93 ± 0.75 69.26 ± 0.10 66.71 ± 0.52
+ MosaicMix [8, 37] 61.15 ± 0.19 58.03 ± 0.76 62.73 ± 0.58 61.18 ± 0.57
+ UniformAug [38] 59.18 ± 0.18 56.35 ± 1.53 65.61 ± 0.37 64.35 ± 0.84
+ TrivialAug [39] 60.22 ± 0.17 57.37 ± 1.48 66.54 ± 0.19 65.34 ± 0.56
+ YOCO [40] 57.52 ± 0.16 54.37 ± 1.70 65.16 ± 0.36 63.12 ± 0.48
+ PixMix [41] 62.07 ± 0.23 58.91 ± 1.13 68.72 ± 0.33 66.82 ± 0.28
+ InterAug [42] 60.94 ± 0.27 58.45 ± 1.20 66.93 ± 0.24 66.52 ± 0.71
+ Ours 63.21 ± 0.18 60.27 ± 1.06 69.78 ± 0.54 67.89 ± 0.46
+ Ours + RandRotate - - 70.37 ± 0.13 -
Swin-T Baseline 61.53 ± 0.49 63.60 ± 0.99 71.03 ± 0.47 68.40 ± 0.68
+ RandRotate [33] 62.54 ± 0.43 64.08 ± 0.78 71.62 ± 0.34 68.52 ± 0.95
+ RandShift [33] 62.77 ± 0.30 63.22 ± 0.73 71.21 ± 0.50 68.51 ± 0.64
+ Cutout [5] 63.17 ± 0.36 63.80 ± 0.88 71.29 ± 0.28 68.41 ± 0.82
+ AutoAug [34] 62.80 ± 0.43 63.19 ± 0.94 70.78 ± 0.64 69.04 ± 0.72
+ RandAug [35] 60.54 ± 0.60 62.07 ± 0.95 68.63 ± 0.41 68.91 ± 0.68
+ AugMix [36] 62.55 ± 0.27 63.07 ± 1.01 70.48 ± 0.61 69.06 ± 0.57
+ MosaicMix [8, 37] 58.26 ± 0.31 60.57 ± 0.76 66.48 ± 0.39 67.24 ± 0.72
+ UniformAug [38] 58.48 ± 0.21 60.55 ± 0.91 67.38 ± 0.45 69.06 ± 0.71
+ TrivialAug [39] 59.72 ± 0.27 61.44 ± 1.12 67.63 ± 0.42 68.53 ± 0.69
+ YOCO [40] 57.05 ± 0.32 58.79 ± 1.09 66.26 ± 0.39 69.04 ± 0.59
+ PixMix [41] 58.15 ± 0.31 58.22 ± 0.81 66.90 ± 0.60 68.15 ± 0.32
+ InterAug [42] 60.72 ± 0.20 62.24 ± 0.99 69.35 ± 0.38 67.09 ± 0.84
+ Ours 62.22 ± 0.23 63.64 ± 0.89 71.30 ± 0.35 68.94 ± 0.59
+ Ours + Cutout 63.18 ± 0.30 - - -
+ Ours + RandRotate - 64.11 ± 0.98 71.69 ± 0.41 -
+ Ours + AugMix - - - 69.12 ± 0.50
IV Conclusion
In this letter, we have proposed a novel data augmentation method named NBBOX for remote sensing object detection. This work thoroughly investigates the effect of bounding box transformation with scaling, rotation, and translation for remote sensing object detection. Besides, NBBOX is easy to integrate with modern deep-learning frameworks and more time-efficient than other augmentation methods. In the future, we will extend our method with various AI paradigms.